# StratLearn-z: Improved photo-$z$ estimation from spectroscopic data subject to selection effects

### v1

@@ Abstract
A precise measurement of photometric redshifts (photo-$z$) is crucial for the success of modern photometric galaxy surveys. Machine learning (ML) methods show great promise in this context, but suffer from covariate shift in training sets due to selection bias where interesting sources, e.g., high redshift objects, are underrepresented, and the corresponding ML models exhibit poor generalisation properties. We present an application of the StratLearn method to the estimation of photo-$z$ (StratLearn-z), validating against simulations where we enforce the presence of covariate shift to different degrees. StratLearn is a statistically principled approach which relies on splitting the combined source and target datasets into strata, based on estimated propensity scores. The latter is the probability for an object in the dataset to be in the source set, given its observed covariates. After stratification, two conditional density estimators are fit separately within each stratum, and then combined via a weighted average. We benchmark our results against the GPz algorithm, quantifying the performance of the two algorithms with a set of metrics. Our results show that the StratLearn-z metrics are only marginally affected by the presence of covariate shift, while GPz shows a significant degradation of performance, specifically concerning the photo-$z$ prediction for fainter objects for which there is little training data. In particular, for the strongest covariate shift scenario considered, StratLearn-z yields a reduced fraction of catastrophic errors, a factor of 2 improvement for the RMSE as well as one order of magnitude improvement on the bias. We also assess the quality of the predicted conditional redshift estimates using the probability integral transform (PIT). The PIT distribution obtained from StratLearn-z features far fewer outliers and is symmetric, i.e., the predictions appear to be well-centered around the true redshift value, despite showing a rather conservative estimation of the spread of the conditional redshift distributions. Our julia implementation of the method, StratLearn-z , is publicly available at https://github.com/chiaramoretti/StratLearn-z.
@@ 1 Introduction
The most precise way to estimate redshifts is via spectroscopic observations. These, however, are demanding in terms of observational time, and challenging for faint, high-redshift objects: the sheer number of sources in future surveys will prevent spectroscopic follow-up for the vast majority. A viable and well established alternative is provided by the so-called photometric redshift (photo-$z$, see e.g., Salvato et al. (2019); Brescia et al. (2021); Newman & Gruen (2022) for detailed reviews), which are extracted from broadband flux measurements performed at different wavelengths. This approach, albeit far less accurate than spectroscopy, allows researchers to measure several objects simultaneously through imaging, as well as extending surveys to deeper regions.
@@
Accuracy and precision in the estimation of photo-$z$ are essential for the success of both current and future imaging surveys. In fact, systematic errors in the determination of redshifts from photometric observations can introduce biases into inferred cosmological parameters. Specifically, the lensing analysis relies on the construction of tomographic redshift bins, with minimal overlap between adjacent bins and a precise determination of the mean redshift of each bin (Ma et al., 2006; Amara & Réfrégier, 2007). On the other hand, the clustering signal used in the 3x2pt analysis requires both the mean redshift and the width of the redshift distribution to be known with high accuracy (Tutusaus et al., 2020; Porredon et al., 2022). There are two main approaches for estimating photo-$z$: template-fitting methods and machine learning (ML) based methods. The former are based on matching the observed photometry to catalogs of template galaxy spectra in order to extract the galaxy redshifts; these methods rely on the completeness of such catalogs. Publicly available template-fitting codes include LePhare (Arnouts et al., 1999; Ilbert et al., 2006), BPZ (Benítez, 2000), Hyperz (Bolzonella et al., 2000) and EAZY (Brammer et al., 2008). On the other hand, the advent of ML techniques has opened new ways to improve photo-$z$ estimation, taking advantage of artificial neural networks (Firth et al., 2003; Collister & Lahav, 2004; Graff et al., 2014) random forests (Carliles et al., 2010; Carrasco Kind & Brunner, 2013), self-organizing maps (Carrasco Kind & Brunner, 2014; Masters et al., 2015), advanced ANNs (Sadeh et al., 2016), Gaussian processes (Almosallam et al., 2016a, b), nearest neighbors (Graham et al., 2018), convolutional neural networks (Pasquet et al., 2019; Henghes et al., 2022) and graph neural networks (Tosone et al., 2023).
@@
Indeed, ML methods have successfully been applied to Stage-III surveys: both the Dark Energy Survey (DES, Abbott et al. (2018)) and the Kilo-Degree Survey (KiDS, Heymans et al. (2021)) have used self-organizing maps to estimate photo-$z$ (see Myles et al. (2021) and Hildebrandt et al. (2021) respectively), while the Hyper Suprime-Cam (HSC, Aihara et al. (2018)) adopted conditional density estimators (Rau et al., 2023; Sugiyama et al., 2023). Performance analyses comparing different approaches have also been presented (Hildebrandt et al., 2010; Sánchez et al., 2014). In preparation for Stage-IV surveys, several ML methods have been compared with each other and to template-fitting methods, in particular for Euclid (Euclid Collaboration et al., 2020) and for LSST-DESC (Schmidt et al., 2020).
@@
ML methods rely on source datasets used to train the algorithms where both spectroscopic and photometric observations are available, in order to learn the relationship between the observed photometry and redshift. A significant issue is posed by non-representative training sets, i.e., spectroscopic training samples that are not random samples from the target photometric survey. Indeed, such non-representative samples are the usual situation in astronomy, where the lack of representativeness is a generic and widespread consequence to selection bias. An example is ‘Malmquist bias’, (Malmquist, 1922, 1925)), which results in spectroscopic catalogs being biased towards brighter, lower redshift sources. Observational biases can thus result in different distributions for the covariates of the source datasets, used to train the ML algorithms, and the target dataset, for which we wish to estimate photo-$z$. Under the assumption that the conditional distribution of the outcome (in this context, the redshift) given the covariates (colors/magnitudes) is the same in the source and target group, the effect is known as covariate shift (Moreno-Torres et al., 2012) in the ML literature. The impact of covariate shift on photo-$z$ estimation has been investigated in Freeman et al. (2017), where the authors propose to mitigate the effect of covariate shift with the use of importance weights, and apply the proposed method to data from DR8 of the Sloan Digital Sky Survey. More recently, Stylianou et al. (2022) investigates the performance of GPz (Almosallam et al., 2016a, b), an algorithm based on Gaussian processes, applying it to simulations that feature imperfections in the spectroscopic training set. Additionally, Toribio San Cipriano et al. (2023) studies the impact of incompleteness of the training set in the context of DES, while Moskowitz et al. (2024) suggests a method to augment the training sample to improve photo-$z$ estimation.
@@
The paper is organized into four additional sections: in Sec. 2 we present the problem of covariate shift and review StratLearn. In Sec. 3 we describe the dataset and the procedure we adopt to introduce covariate shift in the dataset. In Sec. 4 we present our results and comment on them in Sec. 5. Throughout, we refer to the implementation of StratLearn adopted here as StratLearn-z .
@@ 2.2 StratLearn
We fit the two models using the labelled spectroscopic source data as a training set, and obtain an estimate of the redshift distribution for each galaxy in the photometric target data. If the source set is representative of the target set, conditional density estimators commonly aim to minimize the generalized risk under the $L^{2}$-loss:
@@
$\hat{p}(z\mid x)=(1-\alpha)\hat{p}_{\rm Series}(z\mid x)+\alpha~{}\hat{p}_{\rm ker-NN}(z\mid x)\,,$ (3)
where $0\leq\alpha\leq 1$ is a weight parameter. This requires a further optimisation to compute $\alpha$ (within each stratum) via minimisation of the generalised risk:
@@ 3 Simulated Dataset
The dataset was constructed from the catalogs described in Schmidt et al. (2020) (specifically, see Sec. 2 of that work and the references therein for a more detailed description of the dataset). Stylianou et al. (2022) uses this dataset to study the impact of systematics in the training dataset on the the GPz algorithm (Almosallam et al., 2016a, b). Particular attention is paid to two possible degradations in the training dataset that might affect the performance of the algorithm: redshift incompleteness and emission line confusion. The redshift incompleteness in particular is similar to what we explore here, we therefore use GPz to benchmark the performance of StratLearn-z. However, we introduce incompleteness (and therefore, lack of representativeness) in the training data via a different prescription, as detailed in Sec. 3.1.
@@ 3.1 Introducing covariate shift
We enforce covariate shift in the dataset by performing rejection sampling on the the $r$-band, to obtain the target data, following Izbicki et al. (2016). This approach is designed to mimic a selection effect that biases the source set towards brighter objects, similar to what happens with real observations. Specifically, we resample the full dataset via rejection sampling, assigning each galaxy to the target set with probability $p(s=0\mid x)=f_{B(\alpha,\beta)}(x_{r})/\max_{x_{r}}f_{B(\alpha,\beta)}(x_{r})$, where $f_{B(\alpha,\beta)}$ is the density of a beta random variable with parameters $\alpha$, $\beta$ and the $r$-band magnitudes, $x_{r}$, are re-scaled to be between 0 and 1. Galaxies not assigned to the target set are assigned to the source data.
@@
  • no CS: data are randomly split between source and target sets;
  • weak CS: $\alpha=4$, $\beta=5$;
  • mild CS: $\alpha=5$, $\beta=6$;
  • strong CS: $\alpha=5$, $\beta=7$.
@@
Figure 1: Source and target normalised distributions of $r$-band magnitudes for the four covariate shift cases, as stated in each panel. Orange histograms represent the source (labelled) datasets, while in blue we plot the target (unlabelled) datasets. Notice how the rejection sampling procedure introduces a shift – to different degrees, based on the chosen parameters (i.e. scenarios) – between the source and target sets.
Figure 2: Source and target normalised redshift distributions for the four covariate shift cases, as obtained after the rejection-sampling procedure described in Sec. 3.1. Orange histograms represent the source datasets, while in blue we plot the target datasets. We plot the true redshift distribution for the target data, which in reality is of course unknown.
The rejection sampling procedure described above results in source and target datasets of very different sizes among the four CS scenarios. To focus on the impact of covariate shift, and remove dependence of the results on the size of the datasets, we want the source sets for all scenarios to have the same size (and similarly for the target sets). Thus, we randomly resample the sets to construct the source datasets, used to train the conditional density estimators, and target datasets, used to compute performance metrics.
Specifically, we randomly sub-sample 20,000 objects for each training set and 63,000 in each target set, except the target dataset in the weak CS scenario was not sub-sampled because rejection sampling yielded only 62,000 objects.
We choose the target sets to be significantly larger than the training sets to reproduce what happens with real data, where labels (spectroscopic redshifts) are only available for a small fraction of the total photometric measurements. For example, Euclid is expected to observe photometry for $\sim 1.5$ billion galaxies, of which only $\sim 25$ million will have redshift measurements obtained from slitless spectroscopy (Euclid Collaboration et al., 2024).
@@ 3.2 StratLearn computation
For each scenario, we estimate the propensity scores using a logistic regression model to predict the binary classification into the source and target data. In particular, we use the $r$-band data and colors obtained from all other available photometric bands as covariates (i.e., $u-g$, $g-r$, $r-i$, $i-z$, $z-y$). We then partition each of the datasets (combining source and target data under each scenario) into five strata, based on the quintiles of the corresponding propensity score distributions. The resulting strata are populated as detailed in Table 1. The no CS case results in approximately equally sized strata: $\sim 4000$ in the source set and $\sim 12500$ in the target set, per stratum. For the datasets with covariate shift there is significantly more data in the first source strata (corresponding to a lower mean redshift), as expected.
Table 1: Balance between strata after propensity score estimation, for the four covariate shift scenarios. We also report the mean redshift of each stratum for both the source and target datasets. A similar value for the mean redshift within strata indicates that the source and target data are balanced, although we note that this diagnostic is not possible in practice with real source/target data, as $z$ is not observed in the target dataset.
@@
Fig. 3 plots the distributions of the $r$-band magnitudes for the source and target sets within each stratum of the strong CS scenario. Stratification achieves approximately balanced distributions within each strata. Additionally, stratification results in a redshift partition, where in each stratum the mean redshift of the source data approximately matches the mean redshift of the target data – an indication that stratification is approximately removing covariate shift (Autenrieth et al., 2024b) within each stratum. Redshift distribution for source and target sets for the strong CS case are shown in Fig. 15 of Appendix B, showcasing once again the effectiveness of StratLearn to improve balance within each stratum.
Figure 3: Distribution of r-band magnitudes for source (orange histograms) and target (blue histograms) sets for the strong CS scenario, after performing stratification based on estimated propensity scores. In the lower right panel, where we plot the full source and target sets, the distributions are significantly different, while in the other panels, representing each a different stratum, we are able to achieve improved balance.

### v2

@@ Abstract
A precise measurement of photometric redshifts (photo-$z$) is crucial for the success of modern photometric galaxy surveys. Machine learning (ML) methods show great promise in this context, but suffer from covariate shift in training sets due to selection bias where interesting sources, e.g., high redshift objects, are underrepresented, and the corresponding ML models exhibit poor generalisation properties. We present an application of the StratLearn method to the estimation of photo-$z$ (StratLearn-z), validating against simulations where we enforce the presence of covariate shift to different degrees. StratLearn is a statistically principled approach which relies on splitting the combined source and target datasets into strata, based on estimated propensity scores. The latter is the probability for an object in the dataset to be in the source set, given its observed covariates. After stratification, two conditional density estimators are fit separately within each stratum, and then combined via a weighted average. We benchmark our results against the GPz algorithm, quantifying the performance of the two algorithms with a set of metrics. Our results show that the StratLearn-z metrics are only marginally affected by the presence of covariate shift, while GPz shows a significant degradation of performance, specifically concerning the photo-$z$ prediction for fainter objects for which there is little training data. In particular, for the strongest covariate shift scenario considered, StratLearn-z yields a reduced fraction of catastrophic errors, a factor of 2 improvement for the RMSE as well as one order of magnitude improvement on the bias. We also assess the quality of the predicted conditional redshift estimates using the probability integral transform (PIT) and the continuous rank probability score (CRPS). The PIT for StratLearn-z indicates that predictions are well-centered around the true redshift value, if conservative in their variance; the CRPS shows marked improvement at high redshifts when compared with GPz. Our julia implementation of the method, StratLearn-z , is publicly available at https://github.com/chiaramoretti/StratLearn-z.
@@ 1 Introduction
The most precise way to estimate redshifts is via spectroscopic observations. These, however, are demanding in terms of observational time, and challenging for faint, high-redshift objects: the sheer number of sources in future surveys will prevent spectroscopic follow-up for the vast majority. A viable and well established alternative is provided by the so-called photometric redshift (photo-$z$, see e.g., Salvato et al. (2019); Brescia et al. (2021); Newman & Gruen (2022) for detailed reviews), which are extracted from broadband flux measurements performed at different wavelengths. While spectroscopy allows for highly accurate redshift measurements for individual objects, the sample obtained is often biased and unrepresentative of the full population, whereas photometric redshifts, despite being less precise for individual objects, allow the measurement of several objects simultaneously.
@@
Accuracy and precision in the estimation of photo-$z$ are essential for the success of both current and future imaging surveys. In fact, systematic errors in the determination of redshifts from photometric observations can introduce biases into inferred cosmological parameters. Specifically, the lensing analysis relies on the construction of tomographic redshift bins, with minimal overlap between adjacent bins and an accurate determination of the mean redshift of each bin (Ma et al., 2006; Amara & Réfrégier, 2007), as well as knowledge on the $n(z)$. Additionally, the clustering signal used in the 3x2pt analysis requires both the mean redshift and the width of the redshift distribution to be known with high accuracy (Tutusaus et al., 2020; Porredon et al., 2022). There are two main approaches for estimating photo-$z$: template-fitting methods and machine learning (ML) based methods. The former are based on matching the observed photometry to catalogs of template galaxy spectra in order to extract the galaxy redshifts; these methods rely on the completeness of such catalogs. Publicly available template-fitting codes include LePhare (Arnouts et al., 1999; Ilbert et al., 2006), BPZ (Benítez, 2000), Hyperz (Bolzonella et al., 2000) and EAZY (Brammer et al., 2008). On the other hand, the advent of ML techniques has opened new ways to improve photo-$z$ estimation, taking advantage of artificial neural networks (Firth et al., 2003; Collister & Lahav, 2004; Graff et al., 2014) random forests (Carliles et al., 2010; Carrasco Kind & Brunner, 2013), advanced ANNs (Sadeh et al., 2016), Gaussian processes (Almosallam et al., 2016a, b), nearest neighbors (Graham et al., 2018), convolutional neural networks (Pasquet et al., 2019; Henghes et al., 2022), graph neural networks (Tosone et al., 2023), and self-organizing maps (Carrasco Kind & Brunner, 2014; Masters et al., 2015). Self-organizing maps project high-dimensional data points onto a two-dimensional grid. During training, the grid nodes (or neurons) are iteratively adjusted to capture the underlying structure of the data. After training, photometric target galaxies are assigned to their closest grid nodes based on a distance metric (typically Euclidean distance), and photometric redshift (photo-$z$) estimates are derived by averaging the redshifts of the training galaxies associated with each node. This approach can be interpreted as a form of stratification of the data space, where each node effectively defines a stratum. In the original implementations by Carrasco Kind & Brunner (2014) and Masters et al. (2015), the grids consist of approximately 1,000 to 11,000 nodes, making SOMz an extreme example of a stratified approach.
@@
ML methods have successfully been applied to Stage-III surveys: both the Dark Energy Survey (DES, Abbott et al. (2018)) and the Kilo-Degree Survey (KiDS, Heymans et al. (2021)) have used self-organizing maps to estimate photo-$z$ (see Myles et al. (2021) and Hildebrandt et al. (2021) respectively), while the Hyper Suprime-Cam (HSC, Aihara et al. (2018)) adopted conditional density estimators (Rau et al., 2023; Sugiyama et al., 2023). Performance analyses comparing different approaches have also been presented (Hildebrandt et al., 2010; Sánchez et al., 2014). In preparation for Stage-IV surveys, several ML methods have been compared with each other and to template-fitting methods, in particular for Euclid (Euclid Collaboration et al., 2020) and for LSST-DESC (Schmidt et al., 2020).
@@
ML methods rely on source datasets used to train the algorithms where both spectroscopic and photometric observations are available, in order to learn the relationship between the observed photometry and redshift. A significant issue is posed by non-representative training sets, i.e., spectroscopic training samples that are not random samples from the target photometric survey. Indeed, such non-representative samples are the usual situation in astronomy, where the lack of representativeness is a generic and widespread consequence of selection bias. An example is ‘Malmquist bias’, (Malmquist, 1922, 1925)), which results in spectroscopic catalogs being biased towards brighter, lower redshift sources. Observational biases can thus result in different distributions for the covariates of the source datasets, used to train the ML algorithms, and the target dataset, for which we wish to estimate photo-$z$. Under the assumption that the conditional distribution of the outcome (in this context, the redshift) given the covariates (colors/magnitudes) is the same in the source and target group, the effect is known as covariate shift (Moreno-Torres et al., 2012) in the ML literature. The impact of covariate shift on photo-$z$ estimation has been investigated in Freeman et al. (2017), where the authors propose to mitigate the effect of covariate shift with the use of importance weights, and apply the proposed method to data from DR8 of the Sloan Digital Sky Survey. More recently, Stylianou et al. (2022) investigates the performance of GPz (Almosallam et al., 2016a, b), an algorithm based on Gaussian processes, applying it to simulations that feature imperfections in the spectroscopic training set. Additionally, Toribio San Cipriano et al. (2023) studies the impact of incompleteness of the training set in the context of DES, while Moskowitz et al. (2024) suggests a method to augment the training sample to improve photo-$z$ estimation.
@@
Throughout, we refer to the implementation of StratLearn adopted here as StratLearn-z .
@@ 2.2 StratLearn
We fit the two models using the labelled spectroscopic source data as a training set, and obtain an estimate of the redshift probability distribution for each galaxy in the photometric target data. If the source set is representative of the target set, conditional density estimators commonly aim to minimize the generalized risk under the $L^{2}$-loss:
@@
$\hat{p}(z\mid x)=(1-\gamma)\hat{p}_{\rm Series}(z\mid x)+\gamma~{}\hat{p}_{\rm ker-NN}(z\mid x)\,,$ (3)
where $0\leq\gamma\leq 1$ is a weight parameter. This requires a further optimisation to compute $\gamma$ (within each stratum) via minimisation of the generalised risk:
@@ 3 Simulated Dataset
The dataset was constructed from the catalogs described in Schmidt et al. (2020) (specifically, see Sec. 2 of that work and the references therein for a more detailed description of the dataset). Stylianou et al. (2022) uses this dataset to study the impact of systematics in the training dataset on the GPz algorithm (Almosallam et al., 2016a, b). Particular attention is paid to two possible degradations in the training dataset that might affect the performance of the algorithm: redshift incompleteness and emission line confusion. The redshift incompleteness in particular is similar to what we explore here, we therefore use GPz to benchmark the performance of StratLearn-z. However, we introduce incompleteness (and therefore, lack of representativeness) in the training data via a different prescription, as detailed in Sec. 3.1.
@@ 3.1 Introducing covariate shift
We enforce covariate shift in the dataset by performing rejection sampling on the $r$-band, to obtain the target data, following Izbicki et al. (2016). This approach is designed to mimic a selection effect that biases the source set towards brighter objects, similar to what happens with real observations. Specifically, we resample the full dataset via rejection sampling, assigning each galaxy to the target set with probability $p(s=0\mid x)=f_{B(\alpha,\beta)}(x_{r})/\max_{x_{r}}f_{B(\alpha,\beta)}(x_{r})$, where $f_{B(\alpha,\beta)}$ is the density of a beta random variable with parameters $\alpha$, $\beta$ and the $r$-band magnitudes, $x_{r}$, are re-scaled to be between 0 and 1. Galaxies not assigned to the target set are assigned to the source data.
@@
  • no covariate shift: data are randomly split between source and target sets;
  • weak covariate shift: $\alpha=4$, $\beta=5$;
  • mild covariate shift: $\alpha=5$, $\beta=6$;
  • strong covariate shift: $\alpha=5$, $\beta=7$.
@@
Figure 1: Source and target normalised distributions of $r$-band magnitudes for the four covariate shift (CS) cases, as stated in each panel. Orange histograms represent the source (labelled) datasets, while in blue we plot the target (unlabelled) datasets. Notice how the rejection sampling procedure introduces a shift – to different degrees, based on the chosen parameters (i.e. scenarios) – between the source and target sets.
Figure 2: Source and target normalised redshift distributions for the four covariate shift (CS) cases, as obtained after the rejection-sampling procedure described in Sec. 3.1. Orange histograms represent the source datasets, while in blue we plot the target datasets. We plot the true redshift distribution for the target data, which in reality is of course unknown.
The rejection sampling procedure described above results in source and target datasets of very different sizes among the four covariate shift scenarios. To focus on the impact of covariate shift, and remove dependence of the results on the size of the datasets, we want the source sets for all scenarios to have the same size (and similarly for the target sets). Thus, we randomly resample the sets to construct the source datasets, used to train the conditional density estimators, and target datasets, used to compute performance metrics.
Specifically, we randomly sub-sample 20,000 objects for each training set and 63,000 in each target set, except the target dataset in the weak covariate shift scenario was not sub-sampled because rejection sampling yielded only 62,000 objects.
We choose the target sets to be significantly larger than the training sets to reproduce what happens with real data, where labels (spectroscopic redshifts) are only available for a small fraction of the total photometric measurements. For example, Euclid is expected to observe photometry for $\sim 1.5$ billion galaxies, and will rely on complementary ground-based spectroscopic samples to calibrate the relation between galaxy colours and redshifts, with significant effort put into targeted spectroscopic observations to span the full galaxy colour space (see Euclid Collaboration et al. (2024) for more details).
@@ 3.2 StratLearn computation
For each scenario, we estimate the propensity scores using a logistic regression model to predict the binary classification into the source and target data. In particular, we use the $r$-band data and colors obtained from all other available photometric bands as covariates (i.e., $u-g$, $g-r$, $r-i$, $i-z$, $z-y$). We scale all covariates to have mean zero and standard deviation one. We then partition each of the datasets (combining source and target data under each scenario) into five strata, based on the quintiles of the corresponding propensity score distributions. The resulting strata are populated as detailed in Table 1. The no covariate shift case results in approximately equally sized strata: $\sim 4000$ in the source set and $\sim 12500$ in the target set, per stratum. For the datasets with covariate shift there is significantly more data in the first source strata (corresponding to a lower mean redshift), as expected.
Table 1: Balance between strata after propensity score estimation, for the four covariate shift (CS) scenarios. We also report the mean redshift of each stratum for both the source and target datasets. A similar value for the mean redshift within strata indicates that the source and target data are balanced, although we note that this diagnostic is not possible in practice with real source/target data, as $z$ is not observed in the target dataset.
@@
Fig. 3 plots the distributions of the $r$-band magnitudes for the source and target sets within each stratum of the strong covariate shift scenario. Stratification achieves approximately balanced distributions within each strata. Additionally, stratification results in a redshift partition, where in each stratum the mean redshift of the source data approximately matches the mean redshift of the target data – an indication that stratification is approximately removing covariate shift (Autenrieth et al., 2024b) within each stratum. Redshift distribution for source and target sets for the strong covariate shift case are shown in Fig. 16 of Appendix B, showcasing once again the effectiveness of StratLearn to improve balance within each stratum.
Figure 3: Distribution of r-band magnitudes for source (orange histograms) and target (blue histograms) sets for the strong covariate shift (CS) scenario, after performing stratification based on estimated propensity scores. In the lower right panel, where we plot the full source and target sets, the distributions are significantly different, while in the other panels, representing each a different stratum, we are able to achieve improved balance.